OpenStudy (anonymous):

Perimeter of the equilateral triangle ABC is 72 and the perimeter of the isosceles triangle DBC with the base BC=24 is 50. Which is AB + BD? Answer 37 31 27 23

OpenStudy (anonymous):

Since ABC is equilateral, then all sides are equal so AB = BC = AC = 24. The isosceles triangle has two sides that are equal BD and CD so 2 times those lengths + 24 = 50 2x + 24 = 50 2x = 26 x = 13 = BD = CD So AB + BD = 24 +13 = 37

OpenStudy (anonymous):

can you help me on my last question?

OpenStudy (anonymous):

Which question?

OpenStudy (anonymous):

hld on

OpenStudy (anonymous):

Perimeter of the rhombus MNPQ is 48 and the perimeter of the parallelogram NPRS, which has a common side with the rhombus, is 60. Which is the difference between two distinct sides of the parallelogram? Answer 10 9 8 6

OpenStudy (anonymous):

Okay, so since it's a rhombus we know all four sides are the same length. So we know that the common side with the parallelogram is equal to 48 /4 = 12. The opposite or facing sides of a parallelogram are equal, so we know that two of the sides as 12 and 12 = 24. That leaves us with 60 - 24 = 36. 36 / 2 sides = 18. The difference between the sides of the parallelogram are then 18 - 12 = 6.

OpenStudy (anonymous):

Area of the square P is 36 square inches and the perimeter of the rectangle R is 144 inches. If a width of the rectangle is two times a length of the square, which is the length of a length of the rectangle? Answer 70 60 45 30

OpenStudy (anonymous):

can u help me on this one to?

OpenStudy (anonymous):

Okay, the area of the square = 36. A = s x s so the side of the square is equal to 6. The width of the rectangle is 2 times the length of the square so the width = 12. 2 x length + 2 x width = perimeter of rectangle 2l + 2(12) = 144 2l + 24 = 144 2l = 120 l = 60

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

No problem. Helpful to know the properties and equations of area and perimeter of geometric figures.